Two jeepney leaves at the same terminal. Jeepney A heads North at a



constant speed of 40 ??/ℎ while jeepney B heads East at a constant



speed of 60 ??/ℎ. How fast is the distance between them changing



after 30 ?????

1
Expert's answer
2022-05-06T03:25:30-0400

"a^2+b^2=c^2"

For 30 min=0.5 hour one goes 40x0.5=20 km, other 60x0.5=30 km. Si distance between them is "D=\sqrt{20^2+30^2}=36" km

"dD\/dt=36km\/0.5 h=72 km\/hour"